OpenStudy (anonymous):

Given that f(x) : 4x+m and f^-1(x) : nx+3/4, find the values of m and n. Help me T.T

OpenStudy (anonymous):

f(x)=4x+m inverse of f(x)=nx+3/4, f(x)= (x-3/4)/n

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

f(x)=x/n-3/4n x/n=4x, n=1/4 -3/4n= -3=m

OpenStudy (anonymous):

tutor give true answer. tat is invers function so if u want to know more this is the link : sorry bukan jebakan betmen lo. http://en.wikipedia.org/wiki/Inverse_function

OpenStudy (anonymous):

f(x)= y if and only if g(y) = x

OpenStudy (anonymous):

how can he gets f(x)=x/n-3/4n?

OpenStudy (anonymous):

g(y)= x y = nx+3/4 ------ y/n-3/4 n = x so attach the x into f(x) = y f(x) = 4 (y/n-3/4n) + m =0

OpenStudy (anonymous):

well value x =1 4y/n-3n + m =0 4/n-3n+m=0 n= 4/3 and m = 8/3

OpenStudy (anonymous):

okay thanks! finally i got it ^^ thanks noras and tutordiffeq